OpenStudy (anonymous):

Can anyone help? A store recieved 200 containers of juice to be sold by April 1. Each container cost the store $0.75 and sold for $1.25. The store signed a contract with the manufacturer in which the manufacturer agreed to a $0.50 refund for every container not sold by April 1. If 50 containers were sold by April 1, how much profit did the store make?

OpenStudy (anonymous):

total cost => 200*(.75) = 150 total money earned from selling => 50*(1.25) = 62.5 total money earned from refund => 150*(.50) = 75 profit = 150-(62.5+75) profit = $12.5

OpenStudy (anonymous):

thanks

OpenStudy (gw2011):

There is no profit but a loss of $12.50. $62.50 (sale) + $75.00 (refund) = $137.50 (received) $137.50 (received) - $150.00 (cost) = -$12.50 which is a loss